# Your Build Scripts Stink: The State of Code Smells in Build Scripts

### v3

@@ 4. Dynamic Rule Engine
Supporting an additional build system entails providing (i) a parser that emits the canonical data schema and (ii) optional smell detector modules. The dispatcher and rule engine remain untouched, offering a clear path for future expansion (e. g. Bazel or SCons). Details of each of the four build types: parser pattern, regular expressions, the ruleset, and detection logic used behind Sniffer are provided [56].
In summary, Sniffer combines tolerant parsing with a deterministic, modular rule engine, delivering reproducible static analysis across heterogeneous build automation ecosystems while remaining straightforward to extend and maintain.
@@ IV-B2 Evaluation Against Oracle Dataset
We constructed the Oracle dataset using a closed coding approach [77], where a rater analyzes material and identifies code smells based on a predefined codebook [56]. The dataset consists of 72 scripts that were manually examined for code smells. Raters applied their knowledge of programming and build scripts to determine whether a particular smell was present in each script. To avoid bias, we did not include any raters as part of the primary code smells identification or development Sniffer. The smell identification tasks were performed through Qualtrics. In each task, a rater determined which of the code smells identified in Section III were present in a given script. We recruited 20 Computer Science graduate students who have some level of familiarity with build scripts to serve as raters. We obtained an institutional review board (IRB) approval for student participation VII. Each rater was compensated with a \$20 gift card for their involvement. The 72 scripts were sampled to include a mix of scripts with and without code smells. These scripts were then distributed among the 20 raters, ensuring that each script was independently reviewed by at least two raters, with no rater reviewing more than eight scripts. We observe agreements on the scripts, with a Cohen’s Kappa of 0.63. According to Artistein and Poeisio’s interpretation [78], the reported agreement is “substantial”. After the construction of the Oracle dataset, we evaluated the Sniffer’s performance by comparing Oracle results with Sniffers using Precision, Recall, and F1 scores.
@@ IV-C Performance Evaluation Results
Sniffer’s performance for Precision, Recall, and F1 scores is presented in Table III. As shown in the table, the Precision, Recall, and F1 values of the developed tool were measured against the manual analysis, which served as the ground truth. Additionally, we evaluated the tool against the oracle’s output by measuring precision and recall across four types of build scripts: Maven, Gradle, CMake, and Make. When compared against manual analysis, Sniffer achieved consistently high performance, with an average precision of 0.89, a recall of 0.88, and an F1-score of 0.883. Among individual build systems, Maven achieved the highest precision (0.92), while CMake showed the strongest recall (0.93) and F1-score (0.920), indicating reasonably high agreement with ground truth. We also measured the confidence intervals. The 95% confidence interval for manual analysis is (0.852–0.914), indicating a narrow range around the mean, reflecting stable performance. In terms of performance evaluation against the Oracle showed slightly lower but comparable results, with an average precision of 0.83, recall of 0.81, and F1-score of 0.819. Maven again achieved the highest precision (0.84), while both Gradle and CMake obtained the top recall values (0.81). The 95% confidence interval for Oracle evaluation (0.807–0.829) further confirms the consistency of these results. Overall, Sniffer demonstrated strong agreement with manual analysis and reliable performance against the Oracle dataset. The narrow confidence intervals for both comparisons highlight the stability of the tool’s results, underscoring Sniffer’s potential as an effective and reliable automated approach for detecting code smells in build scripts.
@@ V Empirical Analysis of Build Scripts
In this section, we provided answer for RQ2. We performed an empirical analysis of code smell in build scripts using Sniffer. We discussed the build script collection method in Section V-A and the findings for RQ2 in Section V-B.
@@ V-A1 Dataset
In this phase of the study, we conducted an empirical study with a large-scale dataset of Maven, Gradle, CMake, and Make scripts. To examine the prevalence of identified code smells and enhance the generalizability of our results, we focused on GitHub repositories, which are widely used by organizations to host prominent open-source software (OSS) projects [79]. In alignment with established research practices [80], we focused on collecting OSS repositories to have diverse and accessible datasets. The data collection process was followed by a set of pre-defined inclusion criteria, detailed as follows:
  • Criteria 1: The repository should contain at least one of the selected types. The Repositories should contain any of the following types pom.xml, build.gradle, build.gradle.kts, CMake.txt, or Makefile.
@@
To avoid redundancy, repositories that have already been processed and recorded in a central Google Sheet in Section III-A2 data collection were skipped. This automated approach allowed us to collect a large and diverse set of build scripts from GitHub, forming a strong basis for our empirical study and subsequent evaluation of our static analysis tool. We answer RQ2 using 5882 scripts collected from 913, 634, 443, and 2887 repositories, respectively, for Maven, Gradle, CMake, and Make. Summary attributes of the collected repositories are listed in Table IV. Since most repositories contained only a single build script, the total number of build scripts collected is approximately equal to the number of repositories for Maven, Gradle, and Makefiles. Moreover, given the longstanding use of Makefiles in software development, the majority of the collected data consisted of Makefiles.
@@ V-B3 Smell Density
    $\text{Smell Density(d)}=\frac{\text{Total occurrences of x}}{\text{KLOC}}$
@@ VI Discussion
In this section, we recommend mitigation strategies for the identified security smells in the build script and the implementation of Sniffer in Section VI-A, implications of findings VI-B, and provide guidelines for future work  VI-C.
@@ VII Ethical Consideration
This study involved the analysis of publicly available open-source build scripts collected from GitHub. No private or personally identifiable information (PII) was accessed or used. All data was handled in compliance with the platform’s terms of service and policies [94]. We also obtained institutional review board (IRB) approval for our user study. All data was collected, handled, and stored in institutional secure storage. Our goal is to support the broader software engineering community in improving build quality and security.

### v4

@@ 4. Dynamic Rule Engine
Supporting an additional build system entails providing (i) a parser that emits the canonical data schema and (ii) optional smell detector modules. The dispatcher and rule engine remain untouched, offering a clear path for future expansion (e.g., Bazel or SCons). Details of each of the four build types: parser pattern, regular expressions, the ruleset, and detection logic used behind Sniffer are provided [58].
@@ IV-B2 Evaluation Against Oracle Dataset
We constructed the Oracle dataset using a closed coding approach [79], where a set of raters analyzes material and identifies code smells based on a predefined codebook [58]. The material consisted of 72 scripts that were manually examined for code smells. Raters applied their knowledge of programming and build scripts to determine whether a particular smell was present in each script. To avoid bias, raters were not involved in this research as part of the primary code smells identification or the development of Sniffer. As raters, we recruited 20 Computer Science graduate students from NC State University who had some level of familiarity with build scripts to serve as raters. We obtained an institutional review board (IRB) approval for student participation VII. Each rater was compensated with a \$20 gift card for their involvement. The 72 scripts were sampled to include a mix of scripts with and without code smells. These scripts were then distributed among the 20 raters, ensuring that each script was independently reviewed by at least two raters, with no rater reviewing more than eight scripts. The smell identification task was submitted through Qualtrics. In each task, a rater determined which of the code smells identified in Section III were present in a given script. We observe agreements on the scripts, with a Cohen’s Kappa of 0.63. According to Artstein and Poesio’s interpretation [80], the reported agreement is considered “substantial”. After the construction of the Oracle dataset, we evaluated Sniffer’s performance by comparing Oracle results with Sniffer’s findings using Precision, Recall, and F1 scores.
@@ IV-C Performance Evaluation Results
Sniffer’s performance for Precision, Recall, and F1 scores is presented in Table III. As shown in the table, the Precision, Recall, and F1 values of the developed tool were measured against the manual analysis, which served as the ground truth. Additionally, we evaluated the tool against the oracle’s output by measuring precision and recall across four types of build scripts: Maven, Gradle, CMake, and Make. When compared against manual analysis, Sniffer achieved consistently high performance, with an average precision of 0.89, a recall of 0.88, and an F1-score of 0.883. Among individual build systems, Maven achieved the highest precision (0.92), while CMake showed the strongest recall (0.93) and F1-score (0.920), indicating reasonably high agreement with ground truth. We also measured the confidence intervals. The 95% confidence interval for manual analysis is (0.852–0.914), indicating a narrow range around the mean, reflecting stable performance. In terms of performance evaluation against the Oracle, it showed slightly lower but comparable results, with an average precision of 0.83, a recall of 0.81, and an F1-score of 0.819. Maven again achieved the highest precision (0.84), while both Gradle and CMake obtained the top recall values (0.81). The 95% confidence interval for Oracle evaluation (0.807–-0.829) further confirms the consistency of these results. Overall, Sniffer demonstrated strong agreement with manual analysis and reliable performance against the Oracle dataset. The narrow confidence intervals for both comparisons highlight the stability of the tool’s results, underscoring Sniffer’s potential as an effective and reliable automated approach for detecting code smells in build scripts.
@@ V Empirical Analysis of Build Scripts
In this section, we provided an answer for RQ2. We performed an empirical analysis of code smell in build scripts using Sniffer. We discuss the build script collection method in Section V-A and the findings for RQ2 in Section V-B.
@@ V-A1 Dataset
In this phase of the study, we conducted an empirical study with a large-scale dataset of Maven, Gradle, CMake, and Make scripts. To examine the prevalence of identified code smells and enhance the generalizability of our results, we focused on GitHub repositories, which are widely used by organizations to host prominent open-source software (OSS) projects [81]. In alignment with established research practices [82], we focused on collecting OSS repositories to have a diverse and accessible dataset. The data collection process was followed by a set of pre-defined inclusion criteria, detailed as follows:
  • Criteria 1: The repository should contain at least one of the selected types: Maven, Gradle, CMake, Make. The Repositories should contain any of the following types: pom.xml, build.gradle, build.gradle.kts, CMake.txt, or Makefile.
@@
To avoid redundancy, repositories that have already been processed and recorded in a central Google Sheet in Section III-A2 data collection were skipped. This automated approach allowed us to collect a large and diverse set of build scripts from GitHub, forming the basis for the evaluation of the static analysis tool and our empirical study. We answer RQ2 using 5882 scripts collected from 913 (Maven), 634 (Gradle), 443 (CMake), and 2887 (Make) repositories, respectively. Summary attributes of the collected repositories are listed in Table IV. Since most repositories contained only a single build script, the total number of build scripts collected is approximately equal to the number of repositories for Maven, Gradle, and Makefiles. Moreover, given the longstanding use of Makefiles in software development, the majority of the collected data consisted of Makefiles.
@@ V-B3 Smell Density
    $\text{Smell Density(d)}=\frac{\text{Total occurrences of x}}{\text{KLOC}}$ (1)
@@ VI Discussion
In this section, we discuss and recommend mitigation strategies for the identified security smells in the build script in Section VI-A, discuss the implementation of Sniffer for secure build script practice in  VI-B, and provide guidelines for future work in Section VI-C.
@@ VII Ethical Consideration
This study involved the analysis of publicly available open-source build scripts collected from GitHub. No private or personally identifiable information (PII) was accessed or used. All data was handled in compliance with the platform’s terms of service and policies [96]. We also obtained institutional review board (IRB) approval for our user study. All data was collected, handled, and stored in an institutional secure storage. Our goal is to support the broader software engineering community in improving build quality and security.